OpenStudy (anonymous):

Simplify 3 divided by radical 18

OpenStudy (anonymous):

we want to get the radical out of the denominator \[\frac{ 3 }{ \sqrt{18} } = \frac{ 3 }{ \sqrt{9}\sqrt{2} } = \frac{ 3 }{ 3\sqrt{2} } = \frac{ 1 }{ \sqrt{2} } = \frac{ 1 }{ \sqrt{2} }\times \frac{ \sqrt2 }{ \sqrt2 } = \frac{ \sqrt2 }{ 2 }\]

OpenStudy (wolf1728):

SO, you are asking to simplify 3 / sqrt(18) = 3 / sqrt(9*2) = 3 / 3 * sqrt(2) = 1 / sqrt(2)

OpenStudy (anonymous):

@Euler271 This answer is correct.